OpenStudy (richman98):

At which of the following points do the two equations ƒ(x) = 3x2 + 5 and g(x) = 4x + 4 intersect?

OpenStudy (anonymous):

let\[f(x)=g(x)\]

OpenStudy (richman98):

i do not know how to do this plz help

OpenStudy (anonymous):

They intersect at (1,8) If you have more questions similar to this you can go onto desmos.com.

OpenStudy (anonymous):

\[3x ^{2}+5=4x+4\]

OpenStudy (anonymous):

\[3x ^{2}+5-4x-4=0\] \[3x ^{2}-4x+1=0\]

OpenStudy (anonymous):

now can u find the value of x?